### v3

@@ 6.2 Regret Bounds
Denote by $\bar{\pi}^{N}$ the output policy of Algorithm 2 and denote by $N$ the number of rounds. Then, it holds that
$N\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]\leq O\left(\frac{(|\mathcal{S}|+K)^{3}}{\alpha\cdot\sigma\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log(N)}{N}\right)$
where the parameters $\alpha=\min_{k\in[K]}\{\alpha_{k}\}$, $\Delta=\min\{\delta^{2}_{1},\delta^{2}_{2}\}$ with $\delta_{1}$ given in (18) and $\delta_{2}$ given in (21), $\sigma$ given in (28). Also, for any $k\in[K]$, we have
$N\cdot\alpha_{k}-\sum_{n=1}^{N}\hat{\bm{c}}_{k}^{\top}\mathbb{E}\left[\bm{q}^{n}\right]\leq O\left(\frac{(|\mathcal{S}|+K)^{3}}{\alpha\cdot\sigma\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log(N)}{N}\right).$ (33)
@@
We first note that following Theorem 6.1, as long as $n\geq N_{0}$ (or $2N_{0}$ when $N_{0}$ cannot be expressed as the exponent of $2$), we correctly identify the basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ (with a high probability at least $1-O(1/T)$). Then, the resolving step (23) simply becomes
[…]
@@
    $\mathcal{X}=\{\bm{\alpha}^{\prime}\in\mathbb{R}^{|\mathcal{J}^{*}|}:\alpha^{\prime}_{k}\in[\alpha_{k}-\nu,\alpha_{k}+\nu],\forall k\in\mathcal{J}^{*}\},$ (35)
@@
    $\mathcal{Y}=\{\bm{\mu}^{\prime}\in\mathbb{R}^{|\mathcal{S}|}:\mu^{\prime}_{s}\in[\mu_{s}-\nu,\mu_{s}+\nu],\forall s\in\mathcal{S}\}.$ (36)
@@
    $\tau=\min_{n\in[N]}\{\tilde{\bm{\alpha}}(n)\notin\mathcal{X}\text{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\tilde{\bm{\mu}}(n)\notin\mathcal{Y}\}.$ (37)
    and we derive an upper bound of $\mathbb{E}[N-\tau]$, which scales as $O(\log(N))$.
@@
    By using our upper bound of $\mathbb{E}[N-\tau]$ and combining with Lemma 6.2, we show that the gap over the reward $N\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]$ and the gap over the constraint $N\cdot\alpha_{k}-\sum_{n=1}^{N}\hat{\bm{c}}_{k}^{\top}\bm{q}^{n}$ for $k\in\mathcal{J}^{*c}$ can all be upper bounded at the order of $O(\log(N))$. By further exploiting the structure of the matrix $A^{*}$, we can show that the gap over the constraint $N\cdot\alpha_{k}-\sum_{n=1}^{N}\hat{\bm{c}}_{k}^{\top}\bm{q}^{n}$ for $k\in\mathcal{J}^{*}$ can be bounded at the order of $O(\log(N))$ as well.
@@ 6.3 Sample Complexity Bounds
Denote by $\bm{q}^{1},\dots,\bm{q}^{N}$ the output of Algorithm 2. Then, We define $\bar{\bm{q}}^{N}$ such that $\bar{\bm{q}}^{N}_{\mathcal{I}^{*c}}=0$ and
$\bar{\bm{q}}^{N}_{\mathcal{I}^{*}}=\frac{1}{N}\cdot\sum_{n=1}^{N}\bm{q}^{n}_{\mathcal{I}_{N}}.$
Here, $\mathcal{I}^{*}$ can be replaced by $\mathcal{I}_{N}$, the output of Algorithm 1 with the input sample size $N$. From Theorem 6.1, as long as $N\geq N_{0}$ where $N_{0}$ satisfies the condition (26), we know that $\mathcal{I}_{N}=\mathcal{I}^{*}$ with a high probability. We can construct a policy $\bar{\pi}^{N}$ as follows. For each $(s,a)\in\mathcal{S}\times\mathcal{A}$, it holds
$P(\bar{\pi}^{N}(s)=a)=\left\{\begin{aligned} &\frac{\bar{q}^{N}(s,a)}{\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}(s,a^{\prime})},&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}(s,a^{\prime})>0\\ &1/|\mathcal{A}|,&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}\bar{q}^{N}(s,a^{\prime})=0.\end{aligned}\right.$ (38)
We aim to bound the gap between $V_{r}(\pi^{*},\mu_{1})$ and $V_{r}(\bar{\pi}^{N},\mu_{1})$, as well as $V_{k}(\bar{\pi}^{N},\mu_{1})$ and $\alpha_{k}$ for each $k\in[K]$. Note that in the definition (38), the value of $\bar{\pi}^{N}$ is essentially random, where the randomness comes from the filtration $\mathcal{F}_{N}$ that decides the value of $\bm{q}^{1},\dots,\bm{q}^{N}$. Therefore, in the definition of $V_{r}(\bar{\pi}^{N},\mu_{1})$ and $V_{k}(\bar{\pi}^{N},\mu_{1})$ for each $k\in[K]$, we take expectation over the filtration $\mathcal{F}_{N}$. With the above definitions, the regret bound established in Theorem 6.3 can be transferred into the sample complexity bound of the policy $\bar{\pi}^{N}$ given below. The formal proof is relegated to Section 14.
@@
$O\left(\frac{(|\mathcal{S}|+K)^{3}\cdot|\mathcal{S}|\cdot|\mathcal{A}|}{\alpha^{2}\cdot\xi\cdot\sigma(1-\gamma)\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
where $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}$ and $\bm{q}^{*}$ denotes the optimal soluton to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 (defined in (38)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
Note that one caveat in transferring the bounds in Theorem 6.3 into the sample complexity bound of the policy $\bar{\pi}^{N}$ is that the vector $\bar{\bm{q}}^{N}$ is not an occupancy measure since there is a gap between $B\bar{\bm{q}}^{N}$ and $\bm{\mu}$, though bounded at the order of $O(\log N/N)$. However, one can utilize the structure (22) to show a bound over the gap between $\bar{\bm{q}}^{N}$ and $\bm{q}^{*}$, pointwise. In this way, the gap between their state value functions can be bounded, which induces an additional $O(\log(N))$ term. Therefore, we obtain our final sample complexity bounds, scales as $\frac{O(\log^{2}(1/\varepsilon))}{\varepsilon}$.
@@
For the MDP problems without resource constraints, i.e., $K=0$, with a sample complexity bound of
$O\left(\frac{|\mathcal{S}|^{4}\cdot|\mathcal{A}|}{(1-\gamma)^{4}\cdot\xi\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$ (39)
we obtain a policy $\bar{\pi}^{N}$ from Algorithm 2 (defined in (38)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi}^{N},\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\bar{\pi}^{N},\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K].$
Note that one significant term in our sample complexity bound in Proposition 6.5 is $\xi\cdot\Delta$. Recall that $\xi$ represents the minimal mass of the optimal occupancy measure and $\delta$ represents the change of the objective value of LP (8) if we restrict one non-zero basic variable in an optimal solution to be 00. In this sense, $\xi\cdot\Delta$ can be understood as the gap between the value of the optimal policy and the value of a sub-optimal policy that is restricted from taking the optimal action given a state, corresponding to the instance hardness via corner point as described in Section 3.1. Such a problem-dependent gap shows up broadly in the literature for online decision-making with (logarithmic) problem-dependent guarantees, for example, in the multi-arm-bandit problem (Lai and Robbins 1985) and the reinforcement learning problem (Auer et al. 2008).
In terms of the dependency of our sample complexity bound on other problem parameters such as $|\mathcal{S}|$, $|\mathcal{A}|$, and $1-\gamma$, we compare to the series of work (Sidford et al. 2018, Wainwright 2019, Wang 2020, Agarwal et al. 2020, He et al. 2021), that subsequently achieves a sample complexity bound of $O\left(\frac{|\mathcal{S}|\cdot|\mathcal{A}|}{(1-\gamma)^{3}\cdot\varepsilon^{2}}\right)$, where the dependency over $|\mathcal{S}|$, $|\mathcal{A}|$, and $1-\gamma$ is optimal (Gheshlaghi Azar et al. 2013, Lattimore and Hutter 2012). Our sample complexity bound in (39) has a worse dependency in terms of $|\mathcal{S}|$ and $1-\gamma$. This is because we construct an empirical LP to estimate the value of LP (8) and resolve the linear equation as in (23), where the size of the LP (which is $|\mathcal{S}|$) and the eigenvalues of the matrix $A^{*}$ (which is bounded by $1-\gamma$) will play a part. However, our bound (39) enjoys a better dependency in terms of $\varepsilon$.
@@ 7 Extensions
We describe the generalizations of our approach and results in this section. We first describe the generalization to the finite horizon episodic setting, where we maximize the cumulative reward over a finite horizon without discounting. We then describe the generalization to the online learning setting where we can only get the sample of the state we are currently visiting, instead of the sample of an arbitrary state.
@@ 7.1 Finite Horizon Episodic Setting
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{h=1}^{H}c_{h,k}(s_{h},a_{h})\mid\mu_{1}\right]\leq\beta_{k},\leavevmode\nobreak\ \forall k\in[K].$ (41)
@@
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\pi,\mu_{1})-\beta_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K],$ (42)
@@ 7.1.1 LP Reformulation and Estimates
For the finite horizon episodic problem, the occupancy measure is defined as $q_{h,\pi}(s,a)$ for each state-action pair $(s,a)$, each period $h\in[H]$, and the policy $\pi$. Now $q_{h,\pi}(s,a)$ represents the probability of taking action $a$ at period $h$ with state $s$, under policy $\pi$. Following a series of work (e.g. Zimin and Neu (2013), Bhattacharya and Kharoufeh (2017), Efroni et al. (2020)), we obtain the following linear programming to represent the optimal policy and the optimal occupancy measure.
$V^{\mathrm{Fi}}=\leavevmode\nobreak\ \max$ $\sum_{h\in[H]}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\hat{r}_{h}(s,a)\cdot q_{h}(s,a)$ (43a)
@@
Again, we can rewrite LP (43) into the standard formulation LP (8). Now, the matrix $B\in\mathbb{R}^{|\mathcal{S}|H\times|\mathcal{S}||\mathcal{A}|H}$. For $h=1$ and $s\in\mathcal{S}$, the element of $B$ at the $s$-th row and $(s,a,1)$-th column is $1$, for $a\in\mathcal{A}$. For $h\geq 2$ and $s\in\mathcal{S}$, the element of $B$ at $(h-1)\cdot|\mathcal{S}|+s$-th row and $(s,a^{\prime},h)$-th column is $1$, while the element of $B$ at $sh$-th row and $(s^{\prime},a^{\prime},h-1)$-th column is $-P_{h-1}(s|s^{\prime},a^{\prime})$, for $s^{\prime}\in\mathcal{S}$ and $a^{\prime}\in\mathcal{A}$. The element of $B$ at other rows and columns is set to be 00. The matrix $C\in\mathbb{R}^{K\times|\mathcal{S}||\mathcal{S}|H}$ and the element at $k$-th row and $(s,a,h)$-th column is $\hat{c}_{h,k}(s,a)$, for $k\in[K]$ and for $s\in\mathcal{S}$, $a\in\mathcal{A}$, $h\in[H]$. The vector $\bm{\mu}\in\mathbb{R}^{|\mathcal{S}|H}$ and the first $s$-th element of $\bm{\mu}$ is $\mu_{1}(s)$, for $s\in\mathcal{S}$, and the other elements of $\bm{\mu}$ are 00. Then, in the same way as described in Section 4, we can construct $\bar{V}_{N_{0}}$, $\bar{V}_{\mathcal{I},N_{0}}$, and $\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}$ with the newly specified matrix $B$ and $C$.
We now specify the choice of $\lambda_{0}$ and bound the estimation gap for the finite horizon episodic setting. To this end, for an optimal basis $\mathcal{I}^{\prime}$ and $\mathcal{J}^{\prime}$ of LP (43), we denote by
$A^{\prime}=\left[\begin{aligned} &C(\mathcal{J}^{\prime},\mathcal{I}^{\prime})\\ &B(:,\mathcal{I}^{\prime})\end{aligned}\right].$ (44)
We then denote by $\{\sigma_{1}(A^{\prime}),\dots,\sigma_{|\mathcal{S}|+K^{\prime}}(A^{\prime})\}$ the eigenvalues of a matrix $A^{\prime}$. We define $\sigma^{\prime}$ as
$\sigma^{\prime}=\min\left\{|\sigma_{1}(A^{\prime})|,\dots,|\sigma_{|\mathcal{S}|+K^{\prime}}(A^{\prime})|\right\}.$ (45)
We have the following lemma as the bound of the estimation error.
@@
As long as $\lambda_{N_{0}}=\mathrm{Rad}(N_{0},\varepsilon)\cdot H$, the following inequality
$V\leq\bar{V}_{N_{0}}+H\cdot\mathrm{Rad}(N_{0},\varepsilon)\leq$ $V+H\cdot\mathrm{Rad}(N_{0},\varepsilon)+\mathrm{Rad}(N_{0},\varepsilon)\cdot\frac{2(K+|\mathcal{S}|\cdot H)}{\sigma^{\prime}}$ (46)
$+\mathrm{Rad}^{2}(N_{0},\varepsilon)\cdot\frac{K\cdot|\mathcal{S}|+|\mathcal{S}|^{2}\cdot H}{\sigma^{\prime}}.$
holds with probability at least $1-H\cdot(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\varepsilon$.

### v4

@@ 6.2 Regret Bounds
Denote by $N_{2}$ the number of resolving rounds in Algorithm 3. As long as $N_{2}$ satisfies that
$N_{2}\geq\frac{128m^{8}}{(\sigma^{*})^{2}\cdot(1-\gamma)^{2}}\cdot\log(1/\epsilon)$
where $m=|\mathcal{I}^{*}|=|\mathcal{J}^{*}|$, $\epsilon=$, $\sigma^{*}$ defined in (29), it holds that
$N_{2}\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N_{2}}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]\leq\frac{16m^{4}}{\sigma^{*}}+\frac{64m^{7}}{(\sigma^{*})^{3}}\cdot\log(1/\epsilon)$
Also, for any $k\in[K]$, we have
$N_{2}\cdot\alpha_{k}-\sum_{n=1}^{N_{2}}\hat{\bm{c}}_{k}^{\top}\mathbb{E}\left[\bm{q}^{n}\right]\leq\frac{16m^{4}}{\sigma^{*}}+\frac{64m^{7}}{(\sigma^{*})^{3}}\cdot\log(1/\epsilon)$ (34)
and
$\left\|N_{2}\cdot\bm{\mu}-B\cdot\sum_{n=1}^{N_{2}}\mathbb{E}[\bm{q}^{n}]\right\|_{\infty}\leq\frac{16m^{4}}{\sigma^{*}}+\frac{64m^{7}}{(\sigma^{*})^{3}}\cdot\log(1/\epsilon).$ (35)
@@
We first note that following Theorem 6.1, as long as $N_{1}$ satisfies the condition (27), we correctly identify the basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ with a high probability. Then, the resolving step (24) simply becomes
$\left[\begin{aligned} &\bar{C}^{n}(\mathcal{J}_{1}^{*},\mathcal{I}^{*})\\ &\bar{B}^{n}(\mathcal{J}_{2}^{*},\mathcal{I}^{*})\end{aligned}\right]\cdot\tilde{\bm{q}}^{n}_{\mathcal{I}^{*}}=\left[\begin{aligned} &\frac{\bm{\alpha}^{n}_{\mathcal{J}_{1}^{*}}}{N_{2}-n+1}\\ &\frac{\bm{\mu}_{\mathcal{J}^{*}_{2}}^{n}}{N_{2}-n+1}\end{aligned}\right].$ (36)
@@
    $\mathcal{X}=\{\bm{\alpha}^{\prime}\in\mathbb{R}^{|\mathcal{J}_{1}^{*}|}:\alpha^{\prime}_{k}\in[\alpha_{k}-\nu,\alpha_{k}+\nu],\forall k\in\mathcal{J}_{1}^{*}\},$ (37)
@@
    $\mathcal{Y}=\{\bm{\mu}^{\prime}\in\mathbb{R}^{|\mathcal{J}^{*}_{2}|}:\mu^{\prime}_{s}\in[\mu_{s}-\nu,\mu_{s}+\nu],\forall s\in\mathcal{J}^{*}_{2}\}.$ (38)
@@
    $\tau=\min_{n\in[N]}\{\tilde{\bm{\alpha}}(n)\notin\mathcal{X}\text{\penalty 10000\ or\penalty 10000\ }\tilde{\bm{\mu}}(n)\notin\mathcal{Y}\}.$ (39)
    and we derive an upper bound of $\mathbb{E}[N_{2}-\tau]$, which scales as $O(\log(N_{2}))$.
@@
    By using our upper bound of $\mathbb{E}[N_{2}-\tau]$ and combining with Lemma 6.3, we show that the gap over the reward $N_{2}\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N_{2}}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]$ and the gap over the constraint $N_{2}\cdot\alpha_{k}-\sum_{n=1}^{N_{2}}\hat{\bm{c}}_{k}^{\top}\bm{q}^{n}$ for $k\in\mathcal{J}^{*}_{1}$ and $N_{2}\cdot\bm{\mu}_{\mathcal{J}^{*}_{2}}-\sum_{n=1}^{N_{2}}B(\mathcal{J}^{*}_{2},:)\bm{q}^{n}$ can all be upper bounded at the order of $O(\log(N_{2}))$. By further exploiting the structure of the matrix $A^{*}$, we can show that the gap over the constraint $N_{2}\cdot\alpha_{k}-\sum_{n=1}^{N_{2}}\hat{\bm{c}}_{k}^{\top}\bm{q}^{n}$ for $k\in(\mathcal{J}_{1}^{*})^{c}$, and $N_{2}\cdot\bm{\mu}_{(\mathcal{J}^{*}_{2})^{c}}-\sum_{n=1}^{N_{2}}B((\mathcal{J}^{*}_{2})^{c},:)\bm{q}^{n}$ can be bounded at the order of $O(\log(N))$ as well.
@@ 6.3 Sample Complexity Bounds
Denote by $\bm{q}^{1},\dots,\bm{q}^{N_{2}}$ the output of Algorithm 3. Then, We define $\bar{\bm{q}}^{N_{2}}$ such that $\bar{\bm{q}}^{N_{2}}_{\hat{\mathcal{I}}^{c}}=0$ and
$\bar{\bm{q}}^{N_{2}}_{\hat{\mathcal{I}}}=\frac{1}{N_{2}}\cdot\sum_{n=1}^{N_{2}}\bm{q}^{n}_{\hat{\mathcal{I}}}.$
From Theorem 6.1, as long as $N_{1}$satisfies the condition (27), we know that $\hat{\mathcal{I}}$$\mathcal{I}^{*}$ with a high probability. We can construct a policy $\bar{\pi}$ as follows. For each $(s,a)\in\mathcal{S}\times\mathcal{A}$, it holds
[…]
We aim to bound the gap between $V_{r}(\pi^{*},\mu_{1})$ and $V_{r}(\bar{\pi},\mu_{1})$, as well as $V_{k}(\bar{\pi},\mu_{1})$ and $\alpha_{k}$ for each $k\in[K]$. Note that in the definition (40), the value of $\bar{\pi}$ is essentially random, where the randomness comes from the filtration $\mathcal{H}_{N_{2}}$ that decides the value of $\bm{q}^{1},\dots,\bm{q}^{N_{2}}$. Therefore, in the definition of $V_{r}(\bar{\pi},\mu_{1})$ and $V_{k}(\bar{\pi}^{N_{2}},\mu_{1})$ for each $k\in[K]$, we take expectation over the filtration $\mathcal{F}_{N_{2}}$. With the above definitions, the regret bound established in Theorem 6.4 can be transferred into the sample complexity bound of the policy $\bar{\pi}$ given below. The formal proof is relegated to Section 13.6.
@@
$N=O\left(\frac{|\mathcal{S}||\mathcal{A}|\cdot\log(1/\varepsilon)}{\min\{\delta_{1}^{2},\delta_{2}^{2},\sigma_{0}^{2}\}}+\frac{m^{7}\cdot\log(1/\xi)}{\xi(\sigma^{*})^{3}(1-\gamma)}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
where $\xi=\min_{(s,a)\in\mathcal{I}^{*}}\left\{q^{*}(s,a)\right\}>0$ and $\bm{q}^{*}$ denotes the optimal solution to LP (8) corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, we obtain a policy $\bar{\pi}$ from Algorithm 3 (defined in (40)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi},\mu_{1})\leq\varepsilon\text{\penalty 10000\ \penalty 10000\ and\penalty 10000\ \penalty 10000\ }V_{k}(\bar{\pi},\mu_{1})-\alpha_{k}\leq\varepsilon,\penalty 10000\ \forall k\in[K].$
Note that one caveat in transferring the bounds in Theorem 6.4 into the sample complexity bound of the policy $\bar{\pi}$ is that the vector $\bar{\bm{q}}$ is not an occupancy measure since there is a gap between $B\bar{\bm{q}}^{N_{2}}$ and $\bm{\mu}$, though bounded at the order of $O(\log N_{2}/N_{2})$. However, one can utilize the structure (23) to show a bound over the gap between $\bar{\bm{q}}^{N_{2}}$ and $\bm{q}^{*}$, pointwise. In this way, the gap between their state value functions can be bounded, which induces an additional $O(\log(N_{2}))$ term. Therefore, we obtain our final sample complexity bounds, scaling as $\frac{O(\log^{2}(1/\varepsilon))}{\varepsilon}$.
@@
For the MDP problems without resource constraints, i.e., $K=0$, assuming $\xi=O(1/(|\mathcal{S}||\mathcal{A}|))$, with a sample complexity bound of
$O\left(\frac{|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{4}\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$ (41)
with $\Delta=\min\{\delta_{1}^{2},\delta_{2}^{2},\sigma_{0}^{2}\}$, we obtain a policy $\bar{\pi}$ from Algorithm 3 (defined in (40)) such that
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\bar{\pi},\mu_{1})\leq\varepsilon\text{\penalty 10000\ \penalty 10000\ and\penalty 10000\ \penalty 10000\ }V_{k}(\bar{\pi},\mu_{1})-\alpha_{k}\leq\varepsilon,\penalty 10000\ \forall k\in[K].$
Note that one significant term in our sample complexity bound in Proposition 6.6 is $\Delta$. Recall that $\Delta$ represents the change of the objective value of LP (8) if we restrict one non-zero basic variable in an optimal solution to be 0. In this sense, $\Delta$ can be understood as the gap between the value of the optimal policy and the value of a sub-optimal policy that is restricted from taking the optimal action given a state, corresponding to the instance hardness via corner point as described in Section 3.1. Such a problem-dependent gap shows up broadly in the literature for online decision-making with (logarithmic) problem-dependent guarantees, for example, in the multi-arm-bandit problem (lai1985asymptotically) and the reinforcement learning problem (auer2008near).
In terms of the dependency of our sample complexity bound on other problem parameters such as $|\mathcal{S}|$, $|\mathcal{A}|$, and $1-\gamma$, we compare to the series of work (sidford2018near, wainwright2019variance, wang2020randomized, agarwal2020model, he2021nearly), that subsequently achieves a sample complexity bound of $O\left(\frac{|\mathcal{S}|\cdot|\mathcal{A}|}{(1-\gamma)^{3}\cdot\varepsilon^{2}}\right)$, where the dependency over $|\mathcal{S}|$, $|\mathcal{A}|$, and $1-\gamma$ is optimal (gheshlaghi2013minimax, lattimore2012pac). Our sample complexity bound in (41) has a worse dependency in terms of $1-\gamma$. This is because we construct an empirical LP to estimate the value of LP (8) and resolve the linear equation as in (24), where the size of the LP and the singular values of the matrix $A^{*}$ (which is bounded by $1-\gamma$) will play a part. However, our bound (41) enjoys a better dependency in terms of $\varepsilon$. Also, our bound (41) holds under the assumption that $\xi=O(1/(|\mathcal{S}||\mathcal{A}|))$. In the extreme case where $\xi$ can be extremely small, our bound will become worse, scaling with $1/\xi$.
@@ 7 Extensions
We describe the generalizations of our approach and results in this section. We first describe the generalization to the finite horizon episodic setting, where we maximize the cumulative reward over a finite horizon without discounting. We then describe the generalization to the off-policy and on-policy learning settings, which are beyond the generative model in that we can only get the sample of the state we are currently visiting, instead of the sample of an arbitrary state. We show that our approach can be further developed and obtain a similar $\tilde{O}(1/\epsilon)$ sample complexity.
@@ 7.1 Finite Horizon Episodic Setting
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{h=1}^{H}c_{h,k}(s_{h},a_{h})\mid\mu_{1}\right]\leq\beta_{k},\penalty 10000\ \forall k\in[K].$ (43)
@@
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\leq\varepsilon\text{\penalty 10000\ \penalty 10000\ and\penalty 10000\ \penalty 10000\ }V_{k}(\pi,\mu_{1})-\beta_{k}\leq\varepsilon,\penalty 10000\ \forall k\in[K],$ (44)
@@ 7.1.1 LP Reformulation and Estimates
For the finite horizon episodic problem, the occupancy measure is defined as $q_{h,\pi}(s,a)$ for each state-action pair $(s,a)$, each period $h\in[H]$, and the policy $\pi$. Now $q_{h,\pi}(s,a)$ represents the probability of taking action $a$ at period $h$ with state $s$, under policy $\pi$. Following a series of work (e.g. zimin2013online, bhattacharya2017linear, efroni2020exploration), we obtain the following linear programming to represent the optimal policy and the optimal occupancy measure.
$V^{\mathrm{Fi}}=\penalty 10000\ \max\$$\sum_{h\in[H]}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\hat{r}_{h}(s,a)\cdot q_{h}(s,a)$ (45a)
@@
Again, we can rewrite LP (45) into the standard formulation LP (8). Now, the matrix $B\in\mathbb{R}^{|\mathcal{S}|H\times|\mathcal{S}||\mathcal{A}|H}$. For $h=1$ and $s\in\mathcal{S}$, the element of $B$ at the $s$-th row and $(s,a,1)$-th column is $1$, for $a\in\mathcal{A}$. For $h\geq 2$ and $s\in\mathcal{S}$, the element of $B$ at $(h-1)\cdot|\mathcal{S}|+s$-th row and $(s,a^{\prime},h)$-th column is $1$, while the element of $B$ at $sh$-th row and $(s^{\prime},a^{\prime},h-1)$-th column is $P_{h-1}(s|s^{\prime},a^{\prime})$, for $s^{\prime}\in\mathcal{S}$ and $a^{\prime}\in\mathcal{A}$. The element of $B$ at other rows and columns is set to be 0. The matrix $C\in\mathbb{R}^{K\times|\mathcal{S}||\mathcal{S}|H}$ and the element at $k$-th row and $(s,a,h)$-th column is $\hat{c}_{h,k}(s,a)$, for $k\in[K]$ and for $s\in\mathcal{S}$, $a\in\mathcal{A}$, $h\in[H]$. The vector $\bm{\mu}\in\mathbb{R}^{|\mathcal{S}|H}$ and the first $s$-th element of $\bm{\mu}$ is $\mu_{1}(s)$, for $s\in\mathcal{S}$, and the other elements of $\bm{\mu}$ are 0. Then, in the same way as described in Section 4, we can construct $\bar{V}_{N_{0}}$, $\bar{V}_{\mathcal{I},N_{0}}$, and $\text{Dual}_{\mathcal{J},\mathcal{I},N_{0}}$ with the newly specified matrix $B$ and $C$.
We now specify the choice of $\lambda_{0}$ and bound the estimation gap for the finite horizon episodic setting. To this end, we denote by $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ an optimal basis of LP (45), and we denote by
$A^{*}=\left[\begin{aligned} &C(\mathcal{J}^{*}_{1},\mathcal{I}^{*})\\ &B(\mathcal{J}^{*}_{2},\mathcal{I}^{*})\end{aligned}\right].$ (46)
We then define $\sigma^{*}$ as
$\sigma^{*}=\sigma_{\min}(A^{*})$ (47)
where $\sigma_{\min}(a^{*})$ refers to the smallest singular value of the matrix $A^{*}$, where it holds that $\sigma^{*}>0$. We have the following lemma as the bound of the estimation error.
@@
As long as $\lambda_{N_{0}}=\mathrm{Rad}(N_{0},\epsilon)\cdot H$, the following inequality
$V\leq\bar{V}_{N_{0}}+H\cdot\mathrm{Rad}(N_{0},\epsilon)\leq$ $V+H\cdot\mathrm{Rad}(N_{0},\epsilon)+\mathrm{Rad}(N_{0},\epsilon)\cdot\frac{2(K+|\mathcal{S}|\cdot H)}{\sigma^{*}}$ (48)
$+\mathrm{Rad}^{2}(N_{0},\epsilon)\cdot\frac{K\cdot|\mathcal{S}|+|\mathcal{S}|^{2}\cdot H}{\sigma^{*}}.$
holds with probability at least $1-H\cdot(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\epsilon$.